OpenStudy (mathmath333):

if \(\large\tt \begin{align} \color{black}{f(x)=x^3-4x+p}\end{align}\) and \(\large\tt \begin{align} \color{black}{f(0)}\end{align}\) and \(\large\tt \begin{align} \color{black}{f(1)}\end{align}\) are of oppsoite signs ,then which of this is necessarily true?

OpenStudy (mathmath333):

\((a.)-1<p<2<-----\\ (b.)0<p<3\\ (c.)-2<p<1\\ (d.)-3<p<0\)

OpenStudy (mathmath333):

@kirbykirby

OpenStudy (mathmath333):

@aum , @freckles ,help me friends

OpenStudy (aum):

\[ f(x) = x^3 - 4x + p \\ f(0) = 0 - 0 + p = p \\ f(1) = 1 - 4 + p = p - 3 \\ \text{f(0) and f(1) are of opposite signs. Two possibilities:} \\ p \lt 0 ~~\text{ and }~~p-3 > 0 ~~\text{ or } \\ p \gt 0 ~~\text{ and }~~p-3 < 0 \\ \text{First case is not possible. If p < 0, then p - 3 will also be less than 0.} \\ p \gt 0 ~~\text{ and }~~p-3 < 0 \implies 0 \lt p \lt 3 \]

OpenStudy (mathmath333):

thnk u very much @aum